A sample of helium behaves as an ideal gas as it is heated at constant pressure from 253 K to 368 K. If 20 J of work is done by the gas during this process, what is the mass of the helium sample? The universal gas constant is 8.31451 J/mol · K. Answer in units of g.


1
Expert's answer
2022-04-27T13:33:38-0400
"W=\frac{m}{M}R\Delta T"

"m=\frac{MW}{R\Delta T}=\frac{4.0*20}{8.31451*(368-253)}=0.084\:\rm g"